OpenStudy (rayman5766):

what is the value of y so that the ordered pair (5,y) is a solution to the equation -5x + y = -40?

OpenStudy (s4sensitiveandshy):

you have to find y value when x=5 substitute x for 5 solve for y

OpenStudy (rayman5766):

25

OpenStudy (rayman5766):

negative 25

OpenStudy (s4sensitiveandshy):

how did you get that? post your work

OpenStudy (rayman5766):

well 5*5 is 25 but since the first 6 is negative it would make it negative -25

OpenStudy (s4sensitiveandshy):

the first 6 ? what do you mean by that ?

OpenStudy (rayman5766):

not 6 miss typed

OpenStudy (rayman5766):

5

OpenStudy (s4sensitiveandshy):

\[-5x + y = -40 \] replace x with 5 \[-5(5) + y = -40 \] there is negative sign so -5 times 5 = -25 to isolate y you should move the number -25 to the right side

OpenStudy (rayman5766):

thats what i did

OpenStudy (s4sensitiveandshy):

nice!

OpenStudy (s4sensitiveandshy):

so what's the next step ? :)

OpenStudy (rayman5766):

well thats were get stuck

OpenStudy (s4sensitiveandshy):

\[-25+y=-40\] 25 is negative at left side. how would you cancel it out ?

OpenStudy (rayman5766):

25 +40 = 65

OpenStudy (rayman5766):

but since the 25 is a negative it would be _65

OpenStudy (s4sensitiveandshy):

it's -40 that would stay negative you have to add 25 both sides so you can cancel out -25 from left side.

OpenStudy (rayman5766):

-25 +25 = 0

OpenStudy (s4sensitiveandshy):

right what about right side ? because we should add 25 to both sides

OpenStudy (rayman5766):

-40 +25=15

OpenStudy (rayman5766):

-15

OpenStudy (rayman5766):

correct?

OpenStudy (s4sensitiveandshy):

looks good !

OpenStudy (rayman5766):

so am i right

OpenStudy (rayman5766):

wrong

OpenStudy (s4sensitiveandshy):

yes -15 is correct you can check your answer when x =5 y =-15 substitute x and y with these values \[-5(5)+(-15)=-40\] solve left side both sides are equal? that meas (5,-15) is the solution of that equation